Circular No. 7307

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

HD 209458 G. W. Henry, Tennessee State University; G. Marcy, University of

California at Berkeley; R. P. Butler, Department of Terrestrial Magnetism, Carnegie Institution of Washington; and S. S. Vogt, Lick Observatory, report: “The G0 V star HD 209458 = BD +18 4917 (α = 22^{h}03^{m}10^{s}.78, δ = +18^{o}53′03′′.5, equinox 2000.0; V = 7.65) exhibits sinusoidal velocity variations with a semiamplitude of 81 m/s, indicating the presence of a companion with Msini = 0.63 Jupiter mass and an orbital period of 3.523 days. Photometry reveals a transit ingress at Nov. 8.20 UT with a depth of 0.017 mag, consistent with the transit time predicted from the velocities. Further measurements of transits and velocities would be valuable. The next three predicted times of ingress occur on Nov. 15.263, 18.787, and 22.311 (all times uncertain by 1 hr). If correct, the inferred mass is 0.63 that of Jupiter, and radius 1.6 that of Jupiter, implying a density of 0.21 g/cm^{3}. HD 209458 resides at a distance of 47 pc, with rotational v sin i = 3 km/s, and is chromospherically inactive, with an estimated age of 4.5 Gyr.”

SAX J1752.3–3138 M. Cocchi, A. Bazzano, L. Natalucci, and P. Ubertini, Istituto di

Astrofisica Spaziale, CNR, Rome; R. Cornelisse and E. Kuulkers, Space Research Organization Netherlands (SRON), Utrecht, and Astronomical Institute, Utrecht University; and J. in ’t Zand and J. Heise, SRON, communicate: “A post-facto analysis of monitoring observations of the galacticcenter region with the Wide Field Cameras onboard BeppoSAX has led to the discovery of a new x-ray burster, SAX J1752.3–3138. This source was seen only during a burst on Oct. 2.14 UT; no steady emission was observed in 50 000 s net observation time, the upper limit being 6 mCrab. The burster is localized at α = 17^{h}52^{m}24^{s}, δ = −31^{o}37′.7 (equinox 2000.0; 99percent-confidence error radius 2′.9). No previously known x-ray sources are contained within the burst-position confidence circle. The burst showed a peak intensity of 0.7 Crab (2–26 keV), and its spectrum (integrated over 28 s of data) is consistent with a blackbody emission with kT = 1.8 keV. The time profile of the burst is consistent with fast rise and exponential decay. The above characteristics suggest the burst to be a thermonuclear flash on a neutron star, that is likely located in a low-mass x-ray binary.”

1999 November 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green